OpenStudy (anonymous):

Solve. Show all steps. 2(-w - 8) + 9 > -17

OpenStudy (johnweldon1993):

first...distribute that 2 into the parenthesis

OpenStudy (anonymous):

so it would be -2w?

OpenStudy (johnweldon1993):

that's one part...but you also need to do the 2 times -8 part as well

OpenStudy (anonymous):

ok so its -2w * -8 + 9 > -17

OpenStudy (johnweldon1993):

here....distribute...means multiply by each term in the parenthesis..so 2(-w - 8) + 9 > -17 2 times -w = -2w 2 times -8 = -16 so -2w - 16 + 9 > -17 combine -16 + 9 = -7 -2w - 7 > -17 add 7 to both sides -2w > -10 ----------- -2 -2 w < 5 *remember to flip that sign here* hope that is what you got!

OpenStudy (anonymous):

It is thanks Bro